# LASER: Lip Landmark Assisted Speaker Detection for Robustness

### v1

@@ Abstract
Active Speaker Detection (ASD) aims to identify speaking individuals in complex visual scenes. While humans can easily detect speech by matching lip movements to audio, current ASD models struggle to establish this correspondence, often misclassifying non-speaking instances when audio and lip movements are unsynchronized. To address this limitation, we propose Lip landmark Assisted Speaker dEtection for Robustness (LASER). Unlike models that rely solely on facial frames, LASER explicitly focuses on lip movements by integrating lip landmarks in training. Specifically, given a face track, LASER extracts frame-level visual features and the 2D coordinates of lip landmarks using a lightweight detector. These coordinates are encoded into dense feature maps, providing spatial and structural information on lip positions. Recognizing that landmark detectors may sometimes fail under challenging conditions (e.g., low resolution, occlusions, extreme angles), we incorporate an auxiliary consistency loss to align predictions from both lip-aware and face-only features, ensuring reliable performance even when lip data is absent. Extensive experiments across multiple datasets show that LASER outperforms state-of-the-art models, especially in scenarios with desynchronized audio and visuals, demonstrating robust performance in real-world video contexts. Code is available at https://github.com/plnguyen2908/LASER_ASD.
@@ 1 Introduction
Active Speaker Detection (ASD) [10, 26, 25, 15, 19, 2] is a fundamental task in audiovisual computing which aims to detect if one or more people are speaking in a complex visual scene (usually represented by videos). To effectively accomplish this task, models are not only required to understand the visual scene, extract descriptive visual features of candidate speakers, but also learn to correspond the visual information to corresponding audio information. Advanced applications in audiovisual computing such as human-robot interaction [11, 22, 24] and multimodal chatbots [27, 23, 8] usually rely on the success of ASD.
How can we, as humans, tell whether a person is speaking? The synchronization of mouth movements with speech is a natural indicator, allowing us to effortlessly perceive when someone is actually speaking [17]. Even a slight delay that desynchronizes audio and visual cues creates discomfort, making it immediately apparent that the person is not truly speaking at that moment. This inherent sensitivity to audiovisual alignment highlights the central role of mouth movement in active speaker detection.
Figure 1: There is a strong correlation between what a person says and how his/her lips move. As humans, we can easily tell if a person is speaking by corresponding the lip movement and the speech. However, existing state-of-the-art active speaker detection methods (e.g., [10, 26]) do not explicitly encode lip movement, and consequently, struggle to establish this correspondence, often misclassifying non-speaking instances when audio and lip movements are unsynchronized.
Do existing ASD models function in the same intuitive way, and can they detect when a person is not speaking if the audio and mouth movements are unsynchronized? Unfortunately, based on our analysis, the answer to both questions is no. State-of-the-art ASD models consist of two main phases: (1) visual and audio temporal representation learning, which encodes video and audio streams into sequences of visual and audio embeddings with temporal context, and (2) inter- and intra-person context modeling, which leverages cross-attention to capture long-term temporal dependencies and visual-audio interactions. While current approaches often utilize attention mechanisms to model inter- and intra-person context [26, 5, 1, 19, 2], they still struggle to maintain a focus on mouth movement when determining if a person is speaking (Figure 2(a)). This limitation prevents these models from reliably identifying unsynchronization between audio and visual cues, a key factor in human perception.
(a) Grad-CAM Visualization for LoCoNet.
(b) Grad-CAM Visualization for our LASER
@@
Figure 2: Qualitative comparison of Grad-Cam [21] visualizations for LoCoNet and our LASER under nonsynchronized audiovisual scenarios. LoCoNet [26] struggles to accurately predict “not speaking” when visual frames and audio tracks are misaligned, often failing to focus on the lip region when making these predictions. In contrast, our LASER consistently concentrates on the lip area and successfully identifies “not speaking” situations.
@@
While most existing work focuses on improving context modeling, we argue that such limitations stem from the visual features themselves, even before they reach the context modeling module. To address this, we propose Lip landmark Assisted Speaker dEtection for Robustness (LASER) to overcome these limitations. The key novelty of LASER lies in the visual temporal representation learning of ASD models: instead of relying on the model to learn audiovisual interactions purely from facial frames, LASER explicitly directs the model’s attention to lip movements by incorporating lip landmarks in training. By guiding the model to focus specifically on lip landmarks over other facial regions (Figure 2(b)), LASER strengthens the model’s capability to accurately detect active speakers and improves its sensitivity to audiovisual synchronization.
@@
Given a face track [^1] A face track is a sequence of cropped images of the same person’s face across consecutive frames in a video stream [18]., LASER extracts frame-level visual feature maps and the 2D coordinates of lips landmarks for each frame of the face track using a lightweight facial landmark detector [16], forming a “lip track” corresponding it. We then encode the landmark coordinates to continuous 2D feature maps matching the spatial dimension of the visual feature maps and combine them to form our visual features that combine the information from the face track and the lip track. Since the lip landmarks come in as 2D coordinates instead of continuous visual feature maps, integrating such discrete coordinates with continuous feature maps poses a challenge. To address this, we design a simple-yet-effective encoding strategy that converts the 2D discrete coordinate information into continuous 2D visual feature maps matching the spatial dimensions of the corresponding visual features (details in Section 3.2). Furthermore, to alleviate the sparsity of these feature maps, we aggregate them into dense feature maps through a 1x1 convolution layer. This way, the lip landmark coordinates are converted to dense feature maps carrying out the information on the position and form of the lip. The combined visual features are passed through the visual temporal network to obtain the final visual representations with temporal context and the representation can then be used by selected ASD models to capture long-term temporal dependencies and integrate the information from multimodal features.
@@
While LASER enhances ASD model performance, we observe that facial landmark detectors may occasionally fail to provide reliable outputs due to issues like low resolution, occlusion, or extreme facial angles. For example, on widely used AVA benchmarks [18], approximately 15% of face tracks in the test sets lack any landmark detection results. Consequently, missing lip landmarks—resulting in the absence of lip track information—can lead to suboptimal model performance on these test videos. To address this challenge, we introduce an auxiliary consistency loss that aligns predictions made using LASER with those based solely on visual features from the face track. Specifically, for each frame of a face track, we optimize a KL-divergence loss between the predictions obtained with LASER and those based on visual face-frame features alone. This auxiliary loss enables the model to maintain accurate predictions even when the lip track is unavailable, effectively eliminating the dependency on a facial landmark detector at test time.
@@
1) We propose LASER, a novel ASD approach that explicitly directs the model’s attention to lip movements by incorporating lip landmarks in training through effective landmark encoding. 2) We introduce an auxiliary consistency loss during training to address the absence of lip track information at test time, successfully eliminating dependency on facial landmark detectors during inference. 3) Through extensive experiments across multiple datasets and diverse evaluation scenarios, LASER demonstrates stronger performance than state-of-the-art models, particularly in cases where visual frames and audio tracks are unsynchronized, showing robust results in real-world video scenarios.
@@ 2 Related Work
Active Speaker Detection (ASD) remains a crucial challenge in audiovisual computing, aiming to determine if a person is speaking within a given video clip. Earlier works, such as those by Saenko et al. [20] and Everingham et al. [7], concentrated on extracting facial features through key facial region localization. Specifically, Everingham et al. [7] localized and cropped facial regions to compute feature descriptors, while Saenko et al. [20] explicitly localized the lip regions to detect speaking activity. While these prior approaches align with our intuition that speaking behavior is intricately linked to lip movements, they fundamentally differ from our method. Instead of relying exclusively on visual features from the lip region, our approach transforms the 2D coordinates of lip landmarks into sparse feature maps. This technique captures more precise and nuanced details of lip motion, enhancing the performance of ASD.
Recent work in ASD primarily focuses on capturing long-term temporal context [25, 26, 10, 1, 2, 19, 15]. To accomplish this, sequence models such as RNNs [15, 13], attention mechanisms [26, 5], or combinations of both [1] are commonly used to integrate visual and audio features, enabling a more robust understanding of speaking behavior over time. Recently, LoCoNet [26] proposes to use self-attention and cross attention for unimodal and multimodal long-term intra-speaker modeling and further uses CNNs for short term inter-speaker modeling. TalkNCE [10] further advances LoCoNet through an auxiliary multimodal contrastive learning objective. In contrast, our approach shifts the focus in ASD by explicitly guiding models to prioritize lip movements in prediction, providing a complementary enhancement to models centered on contextual modeling. As demonstrated in our experiments, training state-of-the-art models LoCoNet with LASER yields improved performance across various evaluation protocols.
@@ 3 Approach
In this section, we present the details of our proposed Lip landmark Assisted Speaker dEtection for Robustness (LASER). LASER is a versatile training framework that can be integrated with state-of-the-art ASD models. For this work, we adopt the current state-of-the-art model, LoCoNet [26], as our base model to illustrate LASER’s application without loss of generality. We begin by reviewing the LoCoNet framework (Section  3.1), then describe how we incorporate lip movement guidance into the model (Section 3.2), and finally introduce our consistency loss, which addresses the issue of missing lip landmarks at test time (Section 3.3).
@@ 3.1 Preliminaries
State-of-the-art ASD models [26, 25, 15, 5] typically consist of two main components: audio-visual encoders and long-term intra-speaker context modeling modules. The audio-visual encoder includes separate visual and audio encoders, $\mathcal{F}_{v}$ and $\mathcal{F}_{a}$, respectively. The visual encoder $\mathcal{F}_{v}$ processes visual face tracks to produce visual features $f_{v}\in\mathbb{R}^{T\times D}$, while the audio encoder $\mathcal{F}_{a}$ encodes audio spectrograms, generating audio features $f_{a}\in\mathbb{R}^{T\times C}$, where $D$ and $C$ denote the embedding dimensions for visual and audio features, respectively. $f_{v}$ and $f_{a}$ are then concatenated to $f_{av}=concat(f_{v},f_{a})$ to form multimodal features and $f_{av}$ is further processed by the context modeling modules $\mathcal{G}$ [26, 25, 15, 5] to aggregate information from both modalities. Typically, $\mathcal{G}$ outputs multimodal features $f^{\prime}_{av}$ and unimodal features $f^{\prime}_{v}$ and $f^{\prime}_{a}$. Linear classifiers, one for each of the three modalities, are then used to process the corresponding features to make final predictions. $\mathcal{F}$, $\mathcal{G}$ and the linear classifiers are trained end-to-end with the following objective:
@@
where $\mathcal{L}_{av}$ is the cross-entropy loss computed between ground-truth labels $GT$ and predictions $Y$ in each frame and $L_{a}$ and $L_{v}$ are auxiliary cross-entropy losses computed with ground-truth labels and predictions using the unimodal features $f^{\prime}_{v}$ and $f^{\prime}_{a}$, respectively, to avoid $\mathcal{G}$ overly focusing on one modality of features [18].
Figure 3: Illustration of LASER. Given a face track $V$, we first obtain a lip landmark track using a facial landmark detector and encode the 2D coordinates of these landmarks into continuous 2D feature maps. These maps are then aggregated through a 1x1 convolution layer. The encoded lip track is concatenated with visual features from a 3D CNN and fed into ResNet and V-TCN to capture a temporal visual representation, which is further processed by context modeling modules to produce the final prediction. For illustration, we use Long-term Intra-speaker Modeling (LIM) and Short-term Inter-speaker Modeling (SIM) from LoCoNet [26]; however, our LASER is not limited to LoCoNet and can be integrated with other models.
@@ 3.2 Enhanced Representation from Lip Landmark Guidance
While existing work [26, 25, 15, 5] demonstrates competitive performance on the popular AVA-ActiveSpeaker Benchmark [18], our analysis reveals that they still struggles to capture the synchronization between mouth movements and the audio track (Figure 2(a)). Consequently, they often fail to detect non-speaking scenarios when audio and visual frames are misaligned—an inconsistency that humans intuitively recognize. This limitation highlights the need for enhanced focus on lip-audio synchronization within active speaker detection models (Table 4).
@@
We adopt the same visual encoder architecture $\mathcal{F}_{v}$ from prior work [25, 26] where the architecture is composed on a 3D convolution layer, a ResNet-18 [9] backbone, and a video temporal network (V-TCN) [14]. However, unlike prior work which simply forwards visual face tracks to $\mathcal{F}_{v}$, we additionally apply a facial landmark detector to the face track and collect a “lip landmark track” $M$ of the corresponding face track. Each lip landmark is represented as 2D coordinates ${(x_{i},y_{i})}_{i=0}^{K}$ where $K$ is the total number of lip landmarks collected.

### v2

@@ Abstract
Active Speaker Detection (ASD) aims to identify who is speaking in complex visual scenes. While humans naturally rely on lip-audio synchronization, existing ASD models often misclassify non-speaking instances when lip movements and audio are unsynchronized. To address this, we propose Lip landmark Assisted Speaker dEtection for Robustness (LASER), which explicitly incorporates lip landmarks during training to guide the model’s attention to speech-relevant regions. Given a face track, LASER extracts visual features and encodes 2D lip landmarks into dense maps. To handle failure cases such as low resolution or occlusion, we introduce an auxiliary consistency loss that aligns lip-aware and face-only predictions, removing the need for landmark detectors at test time. LASER outperforms state-of-the-art models across both in-domain and out-of-domain benchmarks. To further evaluate robustness in realistic conditions, we introduce LASER-bench, a curated dataset of modern video clips with varying levels of background noise. On the high-noise subset, LASER improves mAP by 3.3 and 4.3 points over LoCoNet and TalkNet, respectively, demonstrating strong resilience to real-world acoustic challenges.
@@ 1 Introduction
Active Speaker Detection (ASD) [jung2024talknce, wang2024loconet, tao2021someone, liao2023light, roy2021learning, alcazar2021maas] is a fundamental task in audiovisual computing which aims to detect if one or more people are speaking in a complex visual scene (usually represented by videos). To effectively accomplish this task, models are not only required to understand the visual scene, extract descriptive visual features of candidate speakers, but also learn to correspond the visual information to corresponding audio information. Advanced applications in audiovisual computing such as human-robot interaction [kang2023video, sheridan2016human, skantze2021turn] and multimodal chatbots [wu2023next, shu2023audio, ge2024worldgpt] usually rely on the success of ASD.
(a) Grad-CAM for LoCoNet.
(b) Grad-CAM for our LASER
@@
Figure 1: Qualitative comparison of Grad-Cam [grad-cam] visualizations for LoCoNet and our LASER under nonsynchronized audiovisual scenarios. LoCoNet [wang2024loconet] struggles to accurately predict “not speaking” when visual frames and audio tracks are misaligned, often failing to focus on the lip region when making these predictions. In contrast, our LASER consistently concentrates on the lip area and successfully identifies “not speaking” situations.
How can we, as humans, tell whether a person is speaking? The synchronization of mouth movements with speech is a natural indicator, allowing us to effortlessly perceive when someone is actually speaking [Park2016LipME]. Even a slight delay that desynchronizes audio and visual cues creates discomfort, making it immediately apparent that the person is not truly speaking at that moment. This inherent sensitivity to audiovisual alignment highlights the central role of mouth movement in active speaker detection.
Do existing ASD models function in the same intuitive way, and can they detect when a person is not speaking if the audio and mouth movements are unsynchronized? Unfortunately, based on our analysis, the answer to both questions is no. State-of-the-art ASD models consist of two main phases: (1) visual and audio temporal representation learning, which encodes video and audio streams into sequences of visual and audio embeddings with temporal context, and (2) inter- and intra-person context modeling, which leverages cross-attention to capture long-term temporal dependencies and visual-audio interactions. While current approaches utilize attention mechanisms to model inter- and intra-person context [wang2024loconet, datta2022asd, alcazar2020active, roy2021learning, alcazar2021maas], they still struggle to maintain a focus on mouth movement when making prediction (Figure 1(a)). This limitation prevents models from reliably identifying unsynchronization between audio and visual cues, a key factor in human perception.
@@
While most existing work focuses on improving context modeling, we argue that such limitations stem from the visual features themselves, even before they reach the context modeling module. To address this, we propose Lip landmark Assisted Speaker dEtection for Robustness (LASER) to overcome these limitations. The key novelty of LASER lies in the lightweight visual temporal representation learning of ASD models: instead of relying on the model to learn audiovisual interactions purely from facial frames, LASER explicitly directs the model’s attention to lip movements by incorporating lip landmarks in training (Figure 1(b)).
@@
Given a face track [^1] A face track is a sequence of cropped images of the same person’s face across consecutive frames in a video stream [roth2020ava]., LASER extracts frame-level visual feature maps and the 2D coordinates of lips landmarks for each frame of the face track using a lightweight facial landmark detector [lugaresi2019mediapipe], forming a “lip track” corresponding it. We then encode the landmark coordinates to continuous 2D feature maps matching the spatial dimension of the visual feature maps and combine them to form our visual features that combine the information from the face track and the lip track. Since the lip landmarks come in as 2D coordinates instead of continuous visual feature maps, integrating such discrete coordinates with continuous feature maps poses a challenge. To address this, we design a simple-yet-effective encoding strategy that converts the 2D discrete coordinate information into continuous 2D visual feature maps matching the spatial dimensions of the corresponding visual features (details in Section 3.2). Furthermore, to alleviate the sparsity of these feature maps, we aggregate them into dense feature maps through a 1x1 convolution layer. This way, the lip landmark coordinates are converted to dense feature maps carrying out the information on the position and form of the lip. The combined visual features are passed through the visual temporal network to obtain the final visual representations with temporal context and the representation can then be used by selected ASD models to capture long-term temporal dependencies and integrate the information from multimodal features.
@@
While LASER enhances ASD model performance, we observe that facial landmark detectors may occasionally fail to provide reliable outputs due to issues like low resolution, occlusion, or extreme facial angles. For example, on widely used AVA benchmarks [roth2020ava], approximately 15% of face tracks in the test sets lack any landmark detection results. Consequently, missing lip landmarks—resulting in the absence of lip track information—can lead to suboptimal model performance on these test videos. To address this challenge, we introduce an auxiliary consistency loss that aligns predictions made using LASER with those based solely on visual features from the face track. Specifically, for each frame of a face track, we optimize a KL-divergence loss between the predictions obtained with LASER and those based on visual face-frame features alone. This auxiliary loss enables the model to maintain accurate predictions even when the lip track is unavailable, eliminating the dependency on a facial landmark detector at test time.
Since LASER leverages explicit lip information, it is particularly suited for acoustically noisy environments where accurate lip-audio correspondence is essential. To evaluate this, we introduce LASER-bench, a curated benchmark featuring modern video content with authentic background noise, including crowd chatter and ambient sounds. Each clip is processed using a voice activity detection (VAD) tool [SileroVAD] to isolate non-speech segments, and categorized into low and high-noise subsets based on RMS energy. This allows for controlled analysis of model robustness under varying noise levels in realistic settings.
To evaluate the effectiveness of our approach, we adopt multiple protocols. Under the standard setup, we train on AVA [roth2020ava] and evaluate on both in-domain (AVA) and out-of-domain (Talkies [alcazar2021maas], ASW [kim2021look]) test sets, where LASER consistently outperforms LoCoNet. To assess noise robustness, we evaluate on LASER-bench, which contains modern video clips grouped into low and high-noise subsets using VAD-filtered RMS thresholds. LASER achieves notable gains over LoCoNet [wang2024loconet] and TalkNet [tao2021someone], particularly under high-noise conditions, up to +3.3 and +4.3 mAP, respectively. We also test robustness to audio-visual misalignment using audio-swapped and delayed audio settings, where our method maintains consistent improvements in frame-level accuracy across all datasets.
@@
1) We propose LASER, a novel ASD approach that explicitly directs the model’s attention to lip movements by incorporating lip landmarks in training through effective landmark encoding. 2) We introduce an auxiliary consistency loss during training to address the absence of lip track information at test time, eliminating dependency on facial landmark detectors during inference. 3) We create LASER-bench, a curated benchmark designed to systematically evaluate ASD robustness under real-world acoustic conditions. 4) LASER outperforms state-of-the-art models under standard evaluation protocols, and further shows improved robustness in real-world scenarios such as audio-visual misalignment and high background noise.
@@ 2 Related Work
Active Speaker Detection.
Active Speaker Detection (ASD) is a key audiovisual challenge, where the goal is to determine if someone is speaking in a video. Recent work [tao2021someone, wang2024loconet, jung2024talknce, alcazar2021maas, roy2021learning, liao2023light] emphasizes long-term temporal modeling via RNNs [liao2023light, kopuklu2021design], attention [wang2024loconet, datta2022asd, ADENet], or both [alcazar2020active]. LoCoNet [wang2024loconet] adopts self- and cross-attention for intra-speaker context, plus CNNs for inter-speaker modeling; TalkNCE [jung2024talknce] extends LoCoNet with contrastive learning. In contrast, our method shifts the focus by explicitly highlighting lip movements, complementing these context-driven approaches. As shown in our experiments, augmenting existing ASD models with LASER yields stronger generalization performance under high background noise context and other complex scenarios.
Facial Landmark Guidance in Audiovisual Computing.
Earlier studies (Saenko et al. [saenko2005visual], Everingham et al. [everingham2009taking]) extracted facial features by localizing key regions. For example, Saenko et al. [saenko2005visual] focused on lip localization to detect speaking activity. Recent works have explored the integration of facial landmarks to improve model performance on audiovisual computing tasks by using facial landmarks for feature pooling [Wang2019landmark_pooling] or directly incorporating discrete landmark coordinates [Geeroms2022_2d_encoding]. While these prior approaches align with our intuition that speaking behavior is intricately linked to lip movements, they fundamentally differ from our method. Our approach transforms the 2D coordinates of lip landmarks into sparse feature maps through a simple-yet-effective encoding function and our approach outperforms these alternative designs within the state-of-the-art approach LoCoNet framework (Section 5.5.1).
Active Speaker Detection Datasets. The field of active speaker detection has been significantly advanced by several key benchmarking datasets. AVA-ActivateSpeaker [roth2020ava] stands out as the largest dataset in this domain, featuring videos derived from movies over a decade ago. In contrast, Talkies [alcazar2021maas] and ASW [kim2021look] curate data from YouTube and the VoxConverse dataset [voxConverse], respectively, focusing on real-world scenarios to provide more diverse speaking environments. Since our approach explicitly incorporates lip information, it is conceptually better equipped to handle acoustically noisy conditions, where strong lip-audio correspondence becomes crucial. To evaluate this, we introduce LASER-bench, a benchmark spanning a wide range of real-world speaking scenarios under background noise for comprehensive assessment.
@@ 3 Approach
In this section, we present the details of our proposed Lip landmark Assisted Speaker dEtection for Robustness (LASER). LASER is a versatile training framework that can be integrated with state-of-the-art ASD models. We begin by reviewing the general framework of ASD models (Section  3.1), then describe how we incorporate lip movement guidance into the model (Section 3.2), and finally introduce our consistency loss, which addresses the issue of missing lip landmarks at test time (Section 3.3).
@@ 3.1 Preliminaries
State-of-the-art ASD models [wang2024loconet, tao2021someone, liao2023light, datta2022asd] typically consist of two main components: audio-visual encoders and long-term intra-speaker context modeling modules. The audio-visual encoder includes separate visual and audio encoders, $\mathcal{F}_{v}$ and $\mathcal{F}_{a}$, respectively. The visual encoder $\mathcal{F}_{v}$ processes visual face tracks to produce visual features $f_{v}\in\mathbb{R}^{T\times D}$, while the audio encoder $\mathcal{F}_{a}$ encodes audio spectrograms, generating audio features $f_{a}\in\mathbb{R}^{T\times C}$, where $D$ and $C$ denote the embedding dimensions for visual and audio features, respectively. $f_{v}$ and $f_{a}$ are then concatenated to $f_{av}=concat(f_{v},f_{a})$ to form multimodal features and $f_{av}$ is further processed by the context modeling modules $\mathcal{G}$ [wang2024loconet, tao2021someone, liao2023light, datta2022asd] to aggregate information from both modalities. Typically, $\mathcal{G}$ outputs multimodal features $f^{\prime}_{av}$ and unimodal features $f^{\prime}_{v}$ and $f^{\prime}_{a}$. Linear classifiers, one for each of the three modalities, are then used to process the corresponding features to make final predictions. $\mathcal{F}$, $\mathcal{G}$ and the linear classifiers are trained end-to-end with the following objective:
@@
where $\mathcal{L}_{av}$ is the cross-entropy loss computed between ground-truth labels $GT$ and predictions $Y$ in each frame and $L_{a}$ and $L_{v}$ are auxiliary cross-entropy losses computed with ground-truth labels and predictions using the unimodal features $f^{\prime}_{v}$ and $f^{\prime}_{a}$, respectively, to avoid $\mathcal{G}$ overly focusing on one modality of features [roth2020ava].
Figure 2: Illustration of LASER. Given a face track $V$, we first obtain a lip landmark track using a facial landmark detector and encode the 2D coordinates of these landmarks into continuous 2D feature maps. These maps are then aggregated through a 1x1 convolution layer. The encoded lip track is concatenated with visual features from a 3D CNN and fed into ResNet and V-TCN to capture a temporal visual representation, which is further processed by context modeling modules [wang2024loconet, tao2021someone, liao2023light] to produce the final prediction. The consistency loss is computed between predictions made with and without lip landmark encoding and gradients are only propagated through the prediction without lip landmark. This way, the model is robust against missing lip landmarks at test time.
@@ 3.2 Enhanced Representation from Lip Landmark Guidance
While existing work [wang2024loconet, tao2021someone, liao2023light, datta2022asd] demonstrates competitive performance on the popular AVA-ActiveSpeaker Benchmark [roth2020ava], our analysis reveals that they still struggles to capture the synchronization between mouth movements and the audio track (Figure 1(a)). This limitation highlights the need for enhanced focus on lip-audio synchronization within active speaker detection models.
@@
We adopt the same visual encoder architecture $\mathcal{F}_{v}$ from prior work [tao2021someone, wang2024loconet] where the architecture is composed on a 3D convolution layer, a ResNet-18 [he2016deep] backbone, and a video temporal network (V-TCN) [lea2016temporal]. However, unlike prior work which simply forwards visual face tracks to $\mathcal{F}_{v}$, we additionally apply a facial landmark detector to the face track and collect a “lip landmark track” $M$ of the corresponding face track. Each lip landmark is represented as 2D coordinates ${(x_{i},y_{i})}_{i=0}^{K}$ where $K$ is the total number of lip landmarks collected.